OpenStudy (anonymous):

simplify: 3-(2x+9)=17

OpenStudy (anonymous):

\(3 - 2x - 9 = 17\) Can you solve this??

OpenStudy (anonymous):

i got that far and then i got 12-2x=17 correct?

OpenStudy (anonymous):

No..

OpenStudy (neo92):

nope.. 3-9 = ?

OpenStudy (anonymous):

\(3 - 9 = ??\)

OpenStudy (anonymous):

i did 3-(-9) *facepalm*

OpenStudy (neo92):

nw u gt it... ??

OpenStudy (anonymous):

\(-6 - 2x = 17\)

OpenStudy (anonymous):

yes, thank you

OpenStudy (anonymous):

Can you further now ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay great..

OpenStudy (anonymous):

Do ask if you have any problem..

OpenStudy (anonymous):

i got the answer -11.5 and i checked it so yippee lol

OpenStudy (anonymous):

Well Done..